The government wants to impose a unit tax on consumers in a market with the following demand and supply: q= 10 - p^d and q=1/2p^s -1/2
a) find the optimal tax rate using the efficiency loss ratio index
b) O= tax revenue - DWL. Using O what is the optimal unit tax rate?
c) Mathematically and on grasp compare your answers of part a and b
d) if government plans to earn exactly the same amount of tax revenue but with AD-Valorem taxon suppliers, what would be the tax rate? show the answer in a generic graph
1
Expert's answer
2020-12-14T05:11:59-0500

a) In equilibrium Qd = Qs:

10 - p = 1/2p - 1/2,

p = 7,

q = 10 - 7 = 3 units.

The tax rate is optimal if the efficiency loss ratio index is minimized, so deadweight loss/tax revenue should be minimized.

b) The tax rate is optimal, if O is minimized too.

c) Answers of part a and b may be different.

d) if government plans to earn exactly the same amount of tax revenue but with AD-Valorem taxon suppliers, then the tax rate should be the same as in a.